# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ Abstract
In higher education, many institutions use algorithmic alerts to flag at-risk students and deliver advising at scale. While much research has focused on evaluating algorithmic predictions, relatively little is known about how discretionary interventions by human experts shape outcomes in algorithm-assisted settings. We study this question using rich quantitative and qualitative data from a randomized controlled trial of an algorithm-assisted advising program at Georgia State University. Taking a mixed-methods approach, we examine whether and how advisors use context unavailable to an algorithm to guide interventions and influence student success. We develop a causal graphical framework for human expertise in the interventional setting, extending prior work on discretion in purely predictive settings. We then test a necessary condition for discretionary expertise using structured advisor logs and student outcomes data, identifying several interventions that meet the criterion for statistical significance. Accordingly, we estimate that 2 out of 3 interventions taken by advisors in the treatment arm were plausibly “expertly targeted” to students using non-algorithmic context. Systematic qualitative analysis of advisor notes corroborates these findings, showing a pattern of advisors incorporating diverse forms of contextual information—such as personal circumstances, financial issues, and student engagement—into their decisions. Our results offer theoretical and practical insight into the real-world effectiveness of algorithm-supported college advising, and underscore the importance of accounting for human expertise in the design, evaluation, and implementation of algorithmic decision systems.
@@
As algorithm–assisted decision making becomes increasingly common in high-stakes public domains, there is growing recognition that evaluating these systems requires more than assessing model performance in isolation. In settings where algorithms inform—but do not replace—human decision-makers, understanding how human experts actually use algorithmic tools is essential to assessing their real-world impact and equity implications [26, 33, 55, 16, 25, 19]. For example, in college advising, an algorithm may identify students at risk of dropping out, but it is up to a human counselor to have a conversation with the student and suggest a course of action that is tailored to the student’s specific circumstances. Despite the widespread deployment of such systems, empirical research remains limited on how human decision-makers interpret algorithmic recommendations, exercise discretion, and influence downstream outcomes.
@@
The Monitoring Advising Analytics to Promote Success (MAAPS) experiment at Georgia State University provides a rare opportunity to investigate these questions in a real-world setting. This multi-year (2016-2020), randomized controlled trial evaluated an algorithm-assisted advising program that offered “intensive, proactive, technology-enhanced advisement and degree planning,” identifying students to support through both advising meetings and algorithmic early warning systems. Notably, the MAAPS program was evaluated via a carefully designed and well-documented experiment that demonstrated positive treatment effects on graduation and persistence [59, 60].
@@
While studies like the MAAPS experiment offer encouraging evidence that algorithmic early warning systems can improve student outcomes, they leave open important questions about why these interventions work. A recent review, for instance, notes that existing studies offer “largely speculative interpretations of intervention mechanisms and effects” [40]. One key question is about the role of human discretion. When algorithms are embedded in human decision processes, it becomes critical to ask: How much of the decision-making success is attributable to the algorithm versus the human expert? What discretionary work do humans perform that the algorithm cannot? What additional information do they need to affect outcomes? These questions help reveal whether non-algorithmic information needs to be explicitly collected by humans, or whether algorithmic systems should be redesigned to better support interventions for student success.
@@
Recognizing that human discretion could shape the impact of algorithm-assisted decision making, our central research question is: What is the evidence for advisor discretion in algorithm-assisted advising, and what is its role in improving student outcomes? Understanding the balance between algorithmic and human contributions is key to scaling algorithm-assisted advising, a strategy that is increasingly used by public universities to boost student success and close equity gaps [57, 21]. While algorithms can be scaled without capacity constraints, the human component cannot — reflecting a hard-won lesson of the social sciences, that interventions performing well in initial studies may not scale due to changes in the implementation, available resources, or personnel training [34, 43]. If human discretion is a key driver of impact, scaling such systems will require not only algorithmic technologies but also investment in human expertise and skills.
@@
To examine the role of human discretion in the success of algorithm-assisted advising, we analyze detailed data on advisor decision-making within the treatment group of the MAAPS trial. The dataset includes structured, time-stamped records of advisor interactions and student outcomes for each semester. In addition to these quantitative logs, MAAPS advisors systematically documented the interventions they delivered during meetings with students (e.g., recommending courses or degree planning) and recorded qualitative comments on their conversations.
@@
Using a novel mixed-methods approach combining theoretical modeling, statistical testing, and qualitative analysis, we investigate one important form of human discretion, how MAAPS advisors use “non-algorithmic information”—information unavailable to an algorithm—to target support to students. In regular meetings, advisors actively gather information about students’ interests, struggles, and personalities, creating an information asymmetry [56] with algorithmic tools that rely on administrative data alone. By leveraging this richer context, advisors may be able to target interventions to students more effectively than would be possible using only the signals provided by any algorithm.
Our study makes two primary contributions:
Contribution 1: Theoretical Model of Expertise
We develop a causal graphical framework to define and audit for human expertise in the interventional setting (§3), adding to a growing body of research studying information asymmetries in human-AI collaboration [7, 28, 31, 4, 5]. Extending prior work on auditing for human expertise in pure prediction tasks [4, 5], we define expertly targeted action in terms of three causal criteria (effectiveness, targeting, and heterogeneity) and analyze a statistical condition that is necessary but insufficient for proving expertise in interventions. We highlight a fundamental impossibility result: expertly targeted action cannot be identified from observational data alone.
Contribution 2: Quantitative and Qualitative Evidence of Expertise
Using quantitative data from the MAAPS trial, we test for expertise by testing the statistical condition analyzed in our theoretical framework, whether logged advisor interventions depend on student outcomes (persistence, GPA) conditional on algorithmically available features. We identify five such potential expert interventions ($p<0.05$) (§4). Subsequent qualitative analysis of advisor comments show that advisors incorporate non-algorithmic context, such as the student’s personal circumstances, career goals, or financial troubles, into these interventions.
Our findings offer evidence for the importance of advisors to the success of data-driven college advising. More broadly, they contribute to a growing body of work examining the role of human discretion in algorithm-assisted interventions.
@@ 1 Related Work
In recent years, there has been a growing interest in the downstream effects of algorithm-assisted human decisions, particularly within criminal justice, economics, healthcare, and higher education [46]. We review key contributions in these areas, contextualizing how they inform our study of algorithm-assisted advising.
Causal Inference in Algorithmic Decision Making.
@@
Causal inference techniques are critical for evaluating algorithm-assisted decisions and their downstream effects. Notably, Imai et al. [33] conducted the first randomized field experiment to evaluate how an algorithmic decision aid in pre-trial risk assessment affects human judgment and decision quality in the U.S. criminal justice system. Follow-up work by Ben-Michael et al. [10] focused on whether algorithm-assisted human decisions improved decision-making “accuracy” over counterfactual baselines. Our work shifts the focus from decision quality per se to the mechanisms of human discretion: we investigate whether, how, and when human decision-makers contribute additional, outcome-relevant expertise. This is especially relevant in the educational setting where advisors choose among multiple possible interventions that may improve student outcomes. Raji and Liu [55] discuss methodological challenges of evaluating algorithmic-assisted decisions arising from the cognitive responses of human decision-makers, emphasizing the need for experimental designs that disentangle the potential cognitive biases of a human judge.
In the educational domain, while the importance of understanding downstream effects of algorithmic student risk assessment has long been recognized [8], there has been little randomized evaluation of algorithmic decision making in education [40]. Perdomo et al. [52] use a regression discontinuity design to estimate the causal effect of early warning system risk scores on student success in Wisconsin public schools, finding small effects but have limited insight into underlying mechanisms due to missing data on subsequent interventions. Liu et al. [44] introduce a causal framework showing that accurate outcome prediction is insufficient for selecting effective interventions to improve individual outcomes in education.
Bias, Expertise, and Discretion in Education and Related Domains.
@@
Discretion can amplify or reduce bias. Green and Chen [26] examine how algorithmic risk scores influence human decision making in a lab study, finding that human evaluators exhibit selective adherence to algorithmic recommendations, often amplifying existing biases. Li [42] further investigates how expertise in peer review influences evaluators’ reliance on algorithmic recommendations, showing that domain experts may either mitigate or exacerbate algorithmic biases depending on their pre-existing heuristics and decision-making frameworks. Likewise, Hoffman et al. [29] analyze how hiring managers integrate algorithmic predictions into their decisions, demonstrating that discretion can either counteract or reinforce algorithmic biases. Gillis et al. [25] show that whether the algorithm amplifies or reduces the bias of the human decision-maker, with respect to a protected class, depends on whether the algorithm uses the protected class as a feature.
Information Asymmetries and Reliance on Algorithms.
Even in settings where algorithms outperform human decision-makers, individuals may possess private information that could refine algorithmic recommendations. Our quantitative approach to studying these information asymmetries is most similar to Alur et al. [4]’s. They devise a statistical test to audit for human use of “algorithmically-unavailable information” in machine-assisted prediction. We apply their test to the MAAPS dataset with slight modifications (described in §4.1). Overall, our work differs significantly with four key contributions. First, we generalize their findings to settings where human experts use non-algorithmic information to take actions rather than make predictions. Second, we formalize the expert decision problem using structural causal modeling and prove identifiability results (§3). Third, we develop a novel mixed-methods analytical framework to complement the quantitative test (§2.2), coding advisor comments to identify the particular non-algorithmic information guiding advisor decision making. Fourth, we apply our methods to a distinct domain—educational interventions in the MAAPS study—whereas Alur et al. [4] study medical triage.
In addition to Alur et al. [4], a growing body of work studies how humans combine their own judgment with the algorithmic predictions and what scenarios lead to humans relying more on algorithms. For example, Balakrishnan et al. [7] show that people often combine their judgments with algorithmic recommendations, increasing prediction errors by over-relying on algorithms when their own insights are more valuable and vice versa. Snyder et al. [67] find that humans rely on algorithms when the algorithms perform better than humans (on average) and when the humans are required to make decisions quickly.
Recent work by Hullman et al. [32] developed a decision-theoretic framework to formalize when human decision-makers can be expected to make utility-maximizing choices in algorithm-assisted contexts, and applied it to quantify the value of information provided by such systems in synthetic decision-making tasks [28]. In contrast, our model of advisor decision making does not assume a rational, utility-maximizing agent, and our goals are to empirically assess whether and how advisors non-algorithmic information in a real world deployment of algorithm-assisted advising.
We discuss further broadly related work in §6.
@@ 2.1.2 Data
For all 1037 students (520 in treatment and 517 in control) who participated in the MAAPS study, the dataset includes their baseline characteristics at the time of enrollment to GSU (such as their demographics, high school GPA and their scores on college entry tests), as well as timeseries data on their educational outcomes updated each term (such as their GPA and credit accumulation), including their graduation outcomes. This data allows us to understand how discretionary decision making connects to students’ educational outcomes.
@@
2.1.3 Data Quality and Cleaning
During the data pre-processing and cleaning phase, we remove students with missing data for transfer status, high school GPA, and expected family contribution (EFC) for all quantitative analyses. [^4] Following exclusion of students missing data on transfer status, all students were non-transfers/first-time freshmen. For all tests run on outcomes derived from timeseries data, we exclude data from summer semesters due to lower student enrollment and different student behaviors over summer semesters as compared to the academic year. Finally, we perform test-specific exclusions, which are described in their respective methods sections.

### v3

@@ Abstract
In higher education, many institutions use algorithmic alerts to flag at-risk students and deliver advising at scale. While much research has focused on evaluating algorithmic predictions, relatively little is known about how discretionary interventions by human experts shape outcomes in algorithm-assisted settings. We study this question using rich quantitative and qualitative data from a randomized controlled trial of an algorithm-assisted advising program at Georgia State University. Taking a mixed-methods approach, we examine whether and how advisors use context unavailable to an algorithm to guide interventions and influence student success. We develop a causal graphical framework for human expertise in the interventional setting, extending prior work on discretion in purely predictive settings. We then test a necessary condition for discretionary expertise using structured advisor logs and student outcomes data, identifying several interventions that meet the criterion for statistical significance. Accordingly, we estimate that 2 out of 3 interventions taken by advisors in the treatment arm were plausibly “expertly targeted” to students using non-algorithmic context. Systematic qualitative analysis of advisor notes corroborates these findings, showing a pattern of advisors incorporating diverse forms of contextual information—such as personal circumstances, financial issues, and student engagement—into their decisions. Finally, we document heterogeneity in advising styles, finding that one style elicits more holistic information about students and is associated with improved graduation rates. Our results offer theoretical and practical insight into the real-world effectiveness of algorithm-supported college advising, and underscore the importance of accounting for human expertise in the design, evaluation, and implementation of algorithmic decision systems.
1 Introduction
@@
As algorithm–assisted decision making becomes increasingly common in high-stakes public domains, there is growing recognition that evaluating these systems requires more than assessing model performance in isolation. In settings where algorithms inform—but do not replace—human decision makers, understanding how human experts actually use algorithmic tools is essential to assessing their real-world impact and equity implications [36, 43, 72, 22, 35, 28]. For example, in college advising, an algorithm may identify students at risk of dropping out, but it is up to a human counselor to have a conversation with the student and suggest a course of action that is tailored to the student’s specific circumstances. Despite the widespread deployment of such systems, empirical research remains limited on how human decision makers interpret algorithmic recommendations, exercise discretion, and influence downstream outcomes.
@@
The Monitoring Advising Analytics to Promote Success (MAAPS) experiment at Georgia State University provides a rare opportunity to investigate these questions in a real-world setting. This multi-year (2016-2022) randomized controlled trial evaluated an algorithm-assisted advising program that offered “intensive, proactive, technology-enhanced advisement and degree planning” [77], identifying students to support through both advising meetings and algorithmic early warning systems. Notably, the MAAPS program was evaluated via a carefully designed and well-documented experiment that demonstrated positive treatment effects on graduation and persistence [76, 77].
@@
While studies like the MAAPS experiment offer encouraging evidence that algorithmic early warning systems can improve student outcomes, they leave open important questions about why these interventions work when they do. A recent review, for instance, notes that existing studies offer “largely speculative interpretations of intervention mechanisms and effects” [50]. One key question is about the role of human discretion. When algorithms are embedded in human-in-the-loop decision processes, it becomes critical to ask: How much of the decision-making success (i.e., improvement in outcomes) is attributable to the algorithm versus the human expert?
@@
Recognizing that human discretion could shape the impact of algorithm-assisted decision making, we hone in on the following research question: What is the evidence for advisor discretion in algorithm-assisted advising, and what is its role in improving student outcomes? Understanding the balance between algorithmic and human contributions is key to scaling algorithm-assisted advising, a strategy that is increasingly used by public universities to boost student success and close equity gaps [74, 30]. While algorithms can be scaled without capacity constraints, the human component cannot—reflecting a hard-won lesson of the social sciences, that interventions performing well in initial studies may not scale due to changes in the implementation, available resources, or personnel training [45, 52, 16]. If human discretion is a key driver of impact, scaling such systems will require not only algorithmic technologies but also investment in human expertise and skills.
@@
To examine the role of human discretion in the success of algorithm-assisted advising, we analyze multifaceted data on advisor decision making within the treatment group of the MAAPS trial. The dataset includes administrative student data that forms the basis for algorithmic alerts, as well as structured, time-stamped records of advisor interactions and student outcomes for each semester (described more thoroughly in §3.1.2). The advisor interaction records include systematic documentation of the interventions they delivered during meetings [^1] Interventions were selected from a list of 20 (see SI A.1.3). with students and qualitative comments on their conversations.
@@
Using a novel mixed-methods approach combining theoretical modeling, statistical testing, and qualitative analysis, we investigate one important form of human discretion, how MAAPS advisors use “non-algorithmic information”—information unavailable to an algorithm—to target support to students. Our study makes two primary contributions:
Contribution 1: Theoretical Model of Expertise.
We develop a causal graphical framework to define and audit for human expertise in the interventional setting (§4), adding to a growing body of research studying information asymmetries in human-AI collaboration [10, 38, 40, 6, 7, 2, 41]. Extending prior work on auditing for human expertise in pure prediction tasks [6, 7], we define expertly targeted action in terms of three causal criteria (effectiveness, targeting, and heterogeneity) and analyze a statistical condition that is necessary but insufficient for proving expertise in interventions. We highlight a fundamental impossibility result: expertly targeted action cannot be identified from observational data alone.
Contribution 2: Quantitative and Qualitative Evidence of Expertise.
We provide empirical evidence for advisor expertise using both quantitative and qualitative data from the MAAPS trial. Quantitatively, we test the statistical condition characterized in our theoretical framework, whether logged advisor interventions depend on student outcomes (persistence, GPA) conditional on algorithmically available features; our analysis is relative to any potential algorithm based on these features. We identify five such potential expert interventions ($p<0.05$) (§5). Qualitative analysis of advisor comments shows that advisors incorporate non-algorithmic context, such as the student’s personal circumstances, career goals, or financial troubles, into these interventions. Finally, we document heterogeneity in advising styles, finding that one style elicits more holistic information about students and is associated with improved graduation rates. Together, these findings emphasize the importance of advisor discretion for administering targeted student support (§6).
Our paper is organized as follows: After a discussion of related work (§2) and our dataset and methodology (§3), §4 introduces a theoretical framework for expertly targeted actions and discusses the implications of testing for expertise using observational data. §5 applies a statistical test of expertise to quantitative data from the MAAPS trial and identifies actions that were plausibly “expertly targeted.” §6 analyzes the qualitative data to further illustrate evidence for expertise. §7 concludes with takeaways for the design, evaluation, and implementation of algorithmic systems.
We recommend readers interested in the mathematical framework include §4. For readers primarily interested in the empirical results, §4 can be skipped.
@@ 2 Related Work
In recent years, there has been a growing interest in the downstream effects of algorithm-assisted human decisions and the importance of human-algorithm collaboration to shaping these impacts. We review key contributions in these areas, contextualizing how they inform our study of algorithm-assisted advising.
2.1 Causal Evaluation of Algorithm-Assisted Decision Making
Predictive vs. Causal Evaluation of Algorithmic Interventions.
@@ 2.1 Causal Evaluation of Algorithm-Assisted Decision Making
Predictive algorithms, such as the student risk assessment tools examined in this study, are typically evaluated by assessing their accuracy against a hold-out dataset. However, a number of factors outside of predictive accuracy can affect the ultimate impact a predictive algorithm has on outcomes: for instance, how well predictions can indicate appropriate interventions [54, 31, 12]. Informed by these limitations of evaluating predictions alone, a growing body of literature instead evaluates algorithms through quantifying the downstream effects of algorithm-assisted human decisions using techniques from causal inference. Notably, Imai et al. [43] conduct the first randomized field experiment to evaluate how an algorithmic decision aid in pretrial risk assessment affects human judgment and decision quality in the U.S. criminal justice system. Follow-up work by Ben-Michael et al. [15] focuses on whether algorithm-assisted human decisions improved decision-making “accuracy” over counterfactual baselines. Raji and Liu [72] discuss methodological challenges of study design given potential cognitive biases of a human judge.
Compared to this prior work, our study shifts the methodological focus from estimating effects of algorithmic interventions per se to the mechanisms of human discretion: we investigate whether, how, and when human decision makers contribute additional, outcome-relevant expertise that enables them to deliver appropriate interventions, improving causal impact.
Causal Evaluation of Learning Analytics Interventions.
In the educational domain, the importance of studying downstream effects of algorithmic student risk assessment is well-recognized [11]. Yet, to the best of our knowledge, there has been little causal evaluation of student risk assessment. For instance, a 2019 review finds that there had been just 11 evaluations of the relationship between any learning analytics interventions and downstream outcomes in higher education, only five of which used randomized designs [50]. Three of these randomized studies examine the effect of using student risk prediction to target interventions like instructor or advisor contact with students, all finding positive impacts on academic performance [44, 56, 63]. There have also been a handful of additional studies aiming to establish causal impact of student risk assessment in US primary and secondary education [27, 58]. In a notable recent example, Perdomo et al. [68] use a regression discontinuity design to estimate the causal effect of intervention targeting using student success risk scores in Wisconsin public schools, finding small effects that they suggest may be equally attainable through school-level interventions.
In general, obstacles to causal evaluation include the logistical difficulties of implementing randomization in educational institutions [25], and collecting longitudinal data on student features, interventions, and outcomes together [11]. Even when randomized or quasi-experimental designs are implemented, we often do not understand why a positive effect occurs due to a lack of qualitative data [50]. The MAAPS study is a rare exception on both fronts: a randomized trial with student features available to a predictive model and detailed structured and qualitative data on advisor decision making, over a period of four years. By studying advisor decision making in the treatment arm of an effective student risk assessment intervention, our work aims to understand the role of human expertise in driving positive impact.
2.2 Human Discretion in Algorithm-Assisted Decision Making
Information Asymmetries
Even in settings where algorithms outperform human decision makers, individuals may possess private information that could refine algorithmic recommendations [62]. Our approach to modeling these information asymmetries is most similar to Alur et al.’s [6]. Alur et al. devise a statistical test to audit for human use of “algorithmically-unavailable information” in improving machine predictions above any possible predictor based on a limited dataset. We apply this test to the MAAPS dataset with slight modifications. Overall, our work differs from Alur et al.’s with four key contributions. First, we generalize Alur et al.’s findings to settings where human experts use non-algorithmic information to take actions rather than make predictions. Second, we formalize the expert decision problem using structural causal modeling and prove identifiability results. Third, we develop a novel mixed-methods analytical framework to complement the quantitative test, coding advisor comments to identify the particular non-algorithmic information guiding advisor decision making. Fourth and finally, while Alur et al. focus on medical prediction, we apply our methods to a distinct domain, educational interventions in the MAAPS study. In this setting, algorithmic context may play a unique “translation” role in bridging from predictions to interventions [55]: Advisors must choose among many possible interventions in order to support students, and non-algorithmic information may help them navigate this expanded action space.
In addition to Alur et al. [6], other work has proposed for studying complementary performance of algorithms and humans in prediction or decision tasks. For example, recent work by Hullman et al. [42] develops a decision-theoretic framework to formalize when human decision makers can be expected to make utility-maximizing choices in algorithm-assisted contexts, which can be applied to quantify the value of information provided by such systems in synthetic decision-making tasks [38]. Additionally, Rastogi et al. [73] develop a general framework in which to express many kinds of machine-human complementarity, including information asymmetries. Donahue et al. [26] also devise a general framework for studying complementary performance. McLaughlin and Spiess [61] devise a causal framework for designing a recommendation algorithm in light of human patterns of compliance to achieve complementary performance. Our study differs from this prior work in its particular focus: Our framework is designed to empirically assess evidence for expertly targeted interventions (rather then predictions, as is the main focus of [73] and [61]), taken by a non-ideal agent in a real-world deployment of an algorithmic aid (rather than a rational, utility-maximizing agent, as with [42]), which improve outcomes through information asymmetries (rather than more general kinds of complementarity, as with [26]).
Reliance
A growing body of work studies how humans combine their own judgment with the algorithmic predictions in practice, by examining what scenarios lead to humans relying more or less on algorithms and whether this improves outcomes. For example, Snyder et al. [85] find that humans rely on algorithms more when the algorithms perform better than humans (on average) and when the humans are required to make decisions quickly. Some work also explicitly examines the role of information asymmetries in reliance. For instance, Balakrishnan et al. [10] show that people may discount prediction errors by over-relying on algorithms when they have access to algorithmically-invisible context that could improve recommendations, Agarwal et al. [2] find that human decision makers make errors by treating AI predictions as statistically independent from their own information, and Hoong et al. [41] find that human decision makers combine their private information with algorithmic predictions more effectively when the predictions are coarsened to lower cognitive burden. Poulidis et al. [70] find evidence that prompting human decision makers with algorithmic attention signals highlighting critical decisions rather than algorithmically-suggested actions may improve decision-making more in the long run, since action suggestions steer decision makers into “uncharted waters” [70]. In contrast to these studies, our goal is not to evaluate how well human agents combine AI predictions and their prior information, but instead to audit for instances of decision makers relying on non-algorithmic information.
2.3 Bias, Expertise and Discretion
@@
Discretion can amplify or reduce bias. For example, Green and Chen [36] examine how algorithmic risk scores influence human decision making in a lab study, finding that human evaluators exhibit selective adherence to algorithmic recommendations, often amplifying existing biases. Li [51] further investigates how expertise in peer review influences evaluators’ reliance on algorithmic recommendations, showing that domain experts may either mitigate or exacerbate algorithmic biases depending on their pre-existing heuristics and decision-making frameworks. Gillis et al. [35] show that whether the algorithm amplifies or reduces the bias of the human decision-maker, with respect to a protected class, depends on whether the algorithm uses the protected class as a feature. We consider the equity impacts of discretionary uptake of algorithmic recommendations in education and beyond in §7.
@@ 3.1.2 Data
For all 1037 students (520 in treatment and 517 in control) who participated in the MAAPS study, the dataset includes their baseline characteristics at the time of enrollment to GSU (such as their demographics, high school GPA and their scores on college entry tests, with a full list provided in Appendix A.1.1), as well as timeseries data on their educational outcomes updated each term (such as their GPA and credit accumulation, with a full list in Appendix A.1.2), including their graduation outcomes. This data allows us to understand how discretionary decision making connects to students’ educational outcomes and to look for heterogeneity of treatment effects in their baseline features.
@@
3.1.3 Data Cleaning
During the data pre-processing and cleaning phase, we remove students with missing data for transfer status and expected family contribution (EFC) for all quantitative analyses. Following exclusion of students missing data on transfer status, all students are non-transfers. Following these exclusions, we have $427$ students in treatment and $423$ in control.
For the $\text{ExpertTest}^{*}$5), we perform additional test-specific exclusions. First, we exclude meeting data from summer semesters due to lower student enrollment and different student behaviors over summer semesters as compared to the academic year. Second, we exclude students missing baseline attribute or outcome data. This results in a test population of $424$ students.